Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for one unit of Zoom are given below:

Standard Quantity or Hours Standard Price
or Rate
Standard
Cost
Direct materials 7.90 pounds $ 2.10 per pound $ 16.59
Direct labor 0.50 hours $ 5.00 per hour $ 2.50

During the most recent month, the following activity was recorded:

  1. Fourteen thousand eight hundred and fifity pounds of material were purchased at a cost of $2.00 per pound.

  2. All of the material purchased was used to produce 1,500 units of Zoom.

  3. 600 hours of direct labor time were recorded at a total labor cost of $4,200.

Required:

1. Compute the materials price and quantity variances for the month.

2. Compute the labor rate and efficiency variances for the month.
Answer #1

Materials Price Variance = (SP-AP) *AQ = (2.10-2)*14,850 = 1485 F Materials Quantity Variance = (SQ-AQ) *SP = (1500*7.90 - 14Labor Rate Variance = (SR-AR) *AH = (5-4200/600)*600 = 1200 U Labor efficiency variance = (SH-AH) *SR = (1500*0.50 - 600)*5 =